OpenStudy (anonymous):

evaluate the expression for a=-1 and b=5. 5a-7

OpenStudy (anonymous):

what do you want?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

a is equal to -1

OpenStudy (anonymous):

i guess it is two

OpenStudy (anonymous):

I guess I kinda forgot how to do this

OpenStudy (anonymous):

5a-7 = 5(-1)-7 = -5-7 = -12